OpenStudy (anonymous):

HELP ASAP ALMOST FINISHED WITH ASSESSMENT THEN DONE WITH DAY OF SCHOOL! Question in comments!

OpenStudy (anonymous):

Solve for K 259/k = 7/4

OpenStudy (amistre64):

what would you suppose for a solution process? this is pretty basic afterall

OpenStudy (anonymous):

cross multiply? idk thats all i got

OpenStudy (amistre64):

thats a good approach, after cross mult, what do we get to work with?

OpenStudy (anonymous):

we have a number we have to subtract from both sides? idk im kinda new to this

OpenStudy (amistre64):

there is no addition or subtraction in the original setup so id go with no. do the cross multiplication first, and lets see what we have to play with

OpenStudy (anonymous):

259 * 4 = 1036